2. Find the price elasticity when p10; p-20 and p-30 for the following demand curve: P 40-1Q 4 a. Remember that the price elasticity formula is: Ed Illustrate the demand below. Now illustrate the price elasticities found above. Do not forget to label your axis b. and points. c. Interpret the price elasticity when the price is 10.
Answer #1

Demand function is P = 40 - Q/4 or Q = 160 - 4P

a) Price elasticity at P = 10 is ed = -4 x 10/(160 - 4*10) = -0.33

Price elasticity at P = 20 is ed = -4 x 20/(160 - 4*20) = -1.00

Price elasticity at P = 30 is ed = -4 x 30/(160 - 4*30) = -3.00

b) Graph is shown below

Demand functiorn 50 40 ed-3.00 30 ed1.00 20 ed--0.33 10 0 0 20 40 60 80100 120 140 160 Quantity

c) When the price is 10, quantity demanded is 120 units. At this price and quantity price elasticity is -0.33 suggesting that a 1 percent increase in price reduces the consumption/quantity demanded by 0.33 percent.